# Breaking through the learning plateaus of in-context learning in Transformer

### v2

@@ Abstract
In-context learning, i.e., learning from context examples, is an impressive ability of Transformer. Training Transformers to possess this in-context learning skill is computationally intensive due to the occurrence of learning plateaus, which are periods within the training process where there is minimal or no enhancement in the model’s in-context learning capability. To study the mechanism behind the learning plateaus, we conceptually seperate a component within the model’s internal representation that is exclusively affected by the model’s weights. We call this the “weights component”, and the remainder is identified as the “context component”. By conducting meticulous and controlled experiments on synthetic tasks, we note that the persistence of learning plateaus correlates with compromised functionality of the weights component. Recognizing the impaired performance of the weights component as a fundamental behavior drives learning plateaus, we have developed three strategies to expedite the learning of Transformers. The effectiveness of these strategies is further confirmed in natural language processing tasks. In conclusion, our research demonstrates the feasibility of cultivating a powerful in-context learning ability within AI systems in an eco-friendly manner.
@@ 1 Introduction
At present, large language models using an autoregressive framework are dominant in the field of machine learning. These large models consume considerable resources to operate, and their size is increasing at a pace that far exceeds the growth of available computing resources [38, 14]. This trend could potentially lead to environmental damage in the future [20, 1, 33, 17]. However, is it essential to rely on massive computational resources and significant energy consumption to develop intelligent systems capable of understanding human language?
@@
Figure 1: A: Examples of the in-context learning tasks. Examples of (1) comes from Alayrac et al. [4], Examples of (2),(3),(4) come from Brown et al. [8]. B: Illustration of learning plateaus and transition pattern. We evaluate the in-context learning ability of Pythia 13B model [6] trained on pile dataset [18] using WordSelection task (Detail in Appendix B) during the training process. During the learning plateaus, there is minimal enhancement in accuracy. However, significant improvement in accuracy is observed during the transition process.
@@
This paper is centered on the in-context learning ability of Transformer, which stands as one of the most significant ability for current applications of Transformer models [8, 15, 39, 31]. Fig. 1A gives the examples of the in-context learning tasks. The in-context learning ability has a confusing property that it is emergent when increasing training FLOPs [47]. By exploring the learning dynamic of in-context learning, previous works  [16, 30, 25, 40, 35] discover that the Transformer learn the in-context learning ability abruptly. We summary the phenomenon during the learning process as plateaus and transition pattern, illustrated in Fig. 1B. This pattern indicates that there is negligible or no enhancement in the in-context learning ability during the initial learning period, which we call the learning plateaus and these plateaus are then followed by rapid and substantial gains, which referred to as the transition process. Often, the strategy to shorten learning plateaus involves expanding the model’s scale, which consequently demands greater computational resources and energy consumption. In this paper, our objective is to investigate the possibility of overcoming the learning plateaus without scaling the model’s size. Addressing this challenge could lead to a novel approach in creating environmentally sustainable intelligent systems. Since learning plateaus are not typically observed in conventional supervised learning, concentrating on the difference may be crucial for unraveling the mechanisms behind the learning plateaus.
@@
Weights and context components  A principal distinction between in-context learning and traditional supervised learning lies in the fact that in-context learning outcomes are shaped by both the model’s parameters and the specific context examples provided. In traditional supervised learning, for a given input sample ${\mathbf{x}}_{p}$ with its corresponding label $y_{p}$, the goal is to define a parameterized function $f_{{\mathbf{w}}}$ with weights ${\mathbf{w}}$, such that the prediction $f_{{\mathbf{w}}}({\mathbf{x}}_{p})$ is equal to the label $y_{p}$. In this scenario, the weights ${\mathbf{w}}$ hold all the information needed to perform the task at hand. Conversely, within the in-context learning paradigm, there is an additional source of information, which is the context examples $s_{c}$. Hence, the prediction model is represented as $f_{{\mathbf{w}},s_{c}}({\mathbf{x}}_{p})$, indicating that both the weights and the context examples have the potential to affect the prediction outcome. To examine how the weights and in-context examples impact the prediction, we assume that the function $f_{{\mathbf{w}},s_{c}}(x)$ can be conceptual decomposite into $f_{{\mathbf{w}},s_{c}}(x_{p})=g_{comb}(g_{weights}({\mathbf{x}}_{p}),g_{context}(s_{c}))$. This decomposition allows us to separate the component $g_{weights}({\mathbf{x}}_{p})$, which is solely dependent on the weights, from the component $g_{context}(s_{c})$, which is influenced by both the weights and the context examples. We refer to $g_{weights}({\mathbf{x}}_{p})$ as the weights component and $g_{context}(s_{c})$ as the context component. Owing to the design of Transformer’s architecture, there is an interaction between $g_{weights}$ and $g_{context}$.
@@
Key observations  Owing to the practical challenges of directly investigating the components, as well as the inability to regulate the complexity of these tasks, we depend on a synthetic task to conduct controlled experiments. Through conducting experiments involving tasks of varying difficulty and monitoring the performance of both the weights and the context elements, we have made a critical observation: the duration of learning plateaus are often associated with a disfunction of the weights component.
@@
Break through the learning plateaus.  Drawing from our observations, we consider the disfunction of the weights component during learning to be the primary cause of the extended learning plateaus. Based on this, we suggest three methods to enhancing the weights component and all these mechods can mitigate the learning plateaus. The effectiveness of these methods are also veried in NLP tasks.
@@
Contributions  Our main contributions can be summarized as follows: a) We give formulations of weights and context components with a new synthetic task that enables the study of the mechanism behind learning plateaus. b) We study the learning process of the synthetic tasks with different complexity. The experiments reveal the relation between the weights component and the learning plateaus. c) We propose methods to break through the learning plateaus from different perspectives and the methods’ effectiveness is also verified in NLP task.
@@ 2 Related works
In this section, we explore the works most closely related to our study. Additional related literature can be found in Appendix C. This appendix encompasses a) an examination of the weights and context components in relation to previously proposed division of in-weights and in-context learning, b) a review of evidence from prior studies which supports the importance of both context and weights components in practical applications, and c) a compilation of related works aimed in understanding. Transformers.
Analyzing the transition phenomenon of in-context learning  The emergence of in-context learning capabilities during learning has captivated numerous researchers [16, 30, 25, 40, 35]. Edelman et al. [16] has found that this transition process is linked to the creation of inductive biases within Transformer models. Meanwhile, other researchers have determined that the occurrence of such a transition is influenced by the properties of the dataset [10]. Li et al. [26], Lu et al. [29] go further to associate the emergent capabilities seen during the learning with those that appear during the scaling phase of model development. Our research, however, is not centered on this transition phenomenon. We aim to delve into the learning plateaus phenomenon, which serves as a vital complement in understanding the intricacies of the transition phenomenon.
Understanding of the mechanism of in-context learning  Considerable research has been devoted to this vital topic, with most prior studies focusing on understanding the mechanism of in-context learning from the perspective of algorithm implementation. For example, a number of recent papers [46, 12, 2] have described in-context learning as akin to performing gradient descent. Additional research [27, 5] has interpreted in-context learning in terms of algorithm implementation and choice. Our study, however, takes a novel approach by examining the in-context learning mechanism through the lens of the weights and contexts components, offering a distinct perspective from earlier works. A more significant difference of our paper is its aim: to overcome the learning plateaus associated with in-context learning.
@@ 3 Experimental Design
This paper employs a synthetic task to investigate the fundamental mechanisms behind learning plateaus. The use of a synthetic task is necessitated by the complexity and unpredictability inherent in real-world tasks. The intricacies of real tasks pose challenges in monitoring and comprehending the precise factors that influence the emergence of learning plateaus.
@@ 3.1 Dataset Construction
We propose a task using the Shapes3D [22] dataset for more controllable study. The experimental setting is shown in Fig. 2. Specifically, given a sequence of image and label pairs as context, the task involves predicting the label of the prompt image. Each image contains six different factors: object color, object shape, object scale, background color, floor color, and pose. We denote the factor as $e$ and the factor value of factor $e$ as $v^{(e)}$. For each sequence, we randomly choose a factor to generate the labels of the images, referring to this factor as the hidden factor $e_{h}$ for this sequence. For the two context sequences in Fig. 2, the hidden factor of Seq #1 is object color, and the correct label for the prompt image is 1 (object color is green). In Seq #2, for the same prompt image, the correct label is 3 (object shape is cube).
@@
Figure 2: Synthetic task. In the task, Transformer is required to predict the label of ${\mathbf{x}}_{p}$ given context examples $s_{c}$. The images from the 3D Shapes dataset are synthesized based on six factors. The output factor is determined by the context. In this case, we provide two sequences of factors: "object color" and "object shape", respectively.
@@ 3.2 Analysis
To successfully tackle the in-context learning task, the network is required to discern the values of the six factors present in the prompt image, which relate to the weight component, as well as accurately determine the appropriate hidden factor for output based on the given contexts, which pertains to the context component. We break down the distribution ${\mathbb{P}}(y_{p}|{\mathbf{x}}_{p},s_{c})$ as follows.
@@
[…]
@@ 3.3 Evaluation framework
In pracetice, since we cannot find the specific form of $g_{weights}$ and $g_{context}$, such that $f_{{\mathbf{w}},s_{c}}(x_{p})=g_{comb}(g_{weights}({\mathbf{x}}_{p}),g_{context}(s_{c}))$. We leverage the probe method to measure the goodness of the components in the inner representation of Transformer as defined in Definition 3.3. We choose the layers in the Transformer that can produce the best probe results. We give the detail of the probe framework in the Appendix A.3.
Probing methods and metrics  We use three metrics here. weights comp. score: accuracy of the probe model to predict $v_{{\mathbf{x}}}$ given ${\mathbf{x}}$, where “comp." is short for component. context comp. score: the accuracy of the probe model to predict $e_{h}$ given $s_{c},{\mathbf{x}}_{p}$. Accuracy: We evaluate the prediction accuracy of the prompt example as the metric to evaluate the in-context learning performance. We give 39 context examples when evaluate context comp. score and the accuracy.
All the results given in this paper are evaluate in the test set by default, as the performance of test set is more aligned with the model’s performance on real situations.
@@ 3.4 Model, training and dataset detail
where $s^{(j)}\triangleq\{({\mathbf{x}}_{1},y_{1}),\cdots,({\mathbf{x}}_{j},y_{j})\}$, $\bm{l}$ denotes the loss function. ${\mathbf{x}}$ will be tokenized by VAE [24] before being passed to Transformer. The training loss in the dataset $S$, which contains $n$ sequence, is calculated as the average of loss over all training sequences, i.e.,
@@
We leave the detail of the model, dataset, the training design and configure in the Appendix A.
@@ 4.1 Controlling of complexity
We aim to understand when the learning plateaus will happen through controlling the complexity of the tasks. Therefore, we will present the method for controlling the complexity in this part. Recall from Definition 3.1 that the data is generated following the formula ${\mathbb{P}}({\mathbf{x}},y,s_{c})=\sum_{m,e_{h}}{\mathbb{P}}(m){\mathbb{P}}(e_{h}){\mathbb{P}}({\mathbf{x}},y,s_{c}|m,e_{h})$. By altering the probabilities ${\mathbb{P}}(m)$, we can manipulate the resulting dataset. We first give two baseline configurations:
@@
When considering two distinct data configuration $D_{1}$ and $D_{2}$, the notation $D_{1}\Rightarrow D_{2}$ denotes the performance evaluation of a model on data setting $D_{2}$ after it has been trained on data setting $D_{1}$. Unless specified otherwise, we assume that the model is both trained and tested on the same data configuration.
@@ 4.2 Learning plateaus of task with different complexity
Learning Plateuas and Transition Point  The primary issue at hand involves the learning plateaus and transition patterns. To begin with, we investigate the capability of the synthetic task to mimic the specific pattern depicted in Fig. 1B. Our scrutiny is directed towards the learning trajectory of the synthetic task. As demonstrated in Fig. 3A, it becomes evident that the synthetic task is successful in replicating the plateaus and transition pattern observed in the actual task.
The length of learning plateus increases when increases the complexity of task  We delved into the relationship between task complexity and the duration of learning plateaus. We employed the entropy of the mapping function $m$’s distribution as an indicator of task complexity. To pinpoint the transition process, we identified the first epoch at which the model achieved a test accuracy greater than 0.17. This threshold was selected because the model’s accuracy remains below 0.17 before reaching the transition process and rises above this thereafter. As anticipated, more complex tasks necessitate longer learning plateaus (Fig. 3B). However, the relationship between the length of the plateaus and the entropy is not linear. With every unit increase in entropy, the extension of the learning plateau is marginal when the entropy is either low or relatively high; conversely, the growth in plateau length is more pronounced at intermediate levels of entropy.
4.3 Disfunction of weights component

### v3

@@ Abstract
In-context learning, i.e., learning from context examples, is an impressive ability of Transformer. Training Transformers to possess this in-context learning skill is computationally intensive due to the occurrence of learning plateaus, which are periods within the training process where there is minimal or no enhancement in the model’s in-context learning capability. To study the mechanism behind the learning plateaus, we conceptually separate a component within the model’s internal representation that is exclusively affected by the model’s weights. We call this the “weights component”, and the remainder is identified as the “context component”. By conducting meticulous and controlled experiments on synthetic tasks, we note that the persistence of learning plateaus correlates with compromised functionality of the weights component. Recognizing the impaired performance of the weights component as a fundamental behavior that drives learning plateaus, we have developed three strategies to expedite the learning of Transformers. The effectiveness of these strategies is further confirmed in natural language processing tasks. In conclusion, our research demonstrates the feasibility of cultivating a powerful in-context learning ability within AI systems in an eco-friendly manner.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: A: Examples of the in-context learning tasks. Examples of (1) comes from Alayrac et al. (2022), Examples of (2),(3),(4) come from Brown et al. (2020). B: Illustration of learning plateaus and transition pattern. We evaluate the in-context learning ability of Pythia 13B model (Biderman et al., 2023) trained on pile dataset (Gao et al., 2020) using WordSelection task (Detail in Appendix B) during the training process.
@@
This paper is centered on the in-context learning ability of Transformer, which stands as one of the most significant abilities for current applications of Transformer models (Brown et al., 2020; Dong et al., 2022; Shin et al., 2022; Min et al., 2021). Fig. 1A gives examples of the in-context learning tasks. The in-context learning ability has a confusing property that it is emergent when increasing training FLOPs (Wei et al., 2022). By exploring the learning dynamic of in-context learning, previous works  (Edelman et al., 2022; Michaud et al., 2023; Kirsch et al., 2022; Singh et al., 2023; Reddy, 2023) discover that the Transformer learns the in-context learning ability abruptly. We summarize the phenomenon during the learning process as plateaus and transition pattern, illustrated in Fig. 1B. This pattern indicates that there is negligible or no enhancement in the in-context learning ability during the initial learning period, which we call the learning plateaus, and these plateaus are then followed by rapid and substantial gains, which are referred to as the transition process. Often, the strategy to shorten learning plateaus involves expanding the model’s scale, which consequently demands greater computational resources and energy consumption. In this paper, our objective is to investigate the possibility of overcoming the learning plateaus without scaling the model’s size. Addressing this challenge could lead to a novel approach to creating environmentally sustainable intelligent systems. Since learning plateaus are not typically observed in conventional supervised learning, concentrating on the difference may be crucial for unraveling the mechanisms behind the learning plateaus.
@@
Weights and context components  A principal distinction between in-context learning and traditional supervised learning lies in the fact that in-context learning outcomes are shaped by both the model’s parameters and the specific context examples provided. In traditional supervised learning, for a given input sample ${\mathbf{x}}_{p}$ with its corresponding label $y_{p}$, the goal is to find a parameterized function $f_{{\mathbf{w}}}$ with weights ${\mathbf{w}}$, such that the prediction $f_{{\mathbf{w}}}({\mathbf{x}}_{p})$ is equal to the label $y_{p}$. In this scenario, the weights ${\mathbf{w}}$ hold all the information needed to perform the task at hand. Conversely, within the in-context learning paradigm, there is an additional source of information, which is the context examples $s_{c}$. Hence, the prediction model is represented as $f_{{\mathbf{w}},s_{c}}({\mathbf{x}}_{p})$, indicating that both the weights and the context examples have the potential to affect the prediction outcome. To examine how the weights and in-context examples impact the prediction, we assume that the function $f_{{\mathbf{w}},s_{c}}(x)$ can be conceptual decomposite into $f_{{\mathbf{w}},s_{c}}(x_{p})=g_{comb}(g_{weights}({\mathbf{x}}_{p}),g_{context}(s_{c}))$. This decomposition allows us to separate the component $g_{weights}(\cdot)$, which is part of the Transformer solely dependent on the weights, from the component $g_{context}(\cdot)$, which is influenced by both the weights and the context examples. We refer to $g_{weights}(\cdot)$ as the weights component and $g_{context}(\cdot)$ as the context component. Owing to the design of the Transformer’s architecture, there is an interaction between $g_{weights}(\cdot)$ and $g_{context}(\cdot)$. Importantly, conceptual decomposition implies that the decomposition isn’t physical; rather, it’s solely for analytical purposes.
@@
Key observations  Owing to the practical challenges of directly investigating the components, as well as the inability to regulate the complexity of these tasks, we depend on a synthetic task to conduct controlled experiments. Through conducting experiments involving tasks of varying difficulty and monitoring the performance of both the weights and the context components, we have made a critical observation: the duration of learning plateaus are often associated with a dysfunction of the weights component.
@@
Break through the learning plateaus.  Drawing from our observations, we consider the dysfunction of the weights component during learning to be the primary cause of the extended learning plateaus. Based on this, we suggest three methods for enhancing the weights component and all these methods can mitigate the learning plateaus. The effectiveness of these methods is also verified in NLP tasks.
@@
Contributions  Our main contributions can be summarized as follows: a) We give formulations of weights and context components with a new synthetic task that enables the study of the mechanism behind learning plateaus. b) We study the learning process of synthetic tasks with different complexity. The experiments reveal the relation between the weights component and the learning plateaus. c) To further verify the causal relation between the learning plateaus and the weights component, we propose different methods to improve the weights component and we observe the mitigating of the learning plateaus. The discoveries are verified in the NLP task.
Figure 2: Synthetic task. In the task, Transformer is required to predict the label of ${\mathbf{x}}_{p}$ given context examples $s_{c}$. The images from the 3D Shapes dataset are synthesized based on six factors. The output factor is determined by the context. In this case, we provide two sequences of factors: ”object color” and ”object shape,” respectively.
@@ 2 Related works
In this section, we explore the works most closely related to our study. Additional related literature can be found in Appendix C. This appendix encompasses a) an examination of the weights and context components in relation to the previously proposed division of in-weights and in-context learning, b) a review of evidence from prior studies that supports the importance of both context and weights components in practical applications, and c) a compilation of related works aimed in understanding Transformers.
Analyzing the transition phenomenon of in-context learning  The emergence of in-context learning capabilities during learning has captivated numerous researchers (Olsson et al., 2022; Michaud et al., 2023; Kirsch et al., 2022; Singh et al., 2023; Reddy, 2023). Olsson et al. (2022) has found that this transition process is linked to the creation of inductive heads within Transformer models. Meanwhile, other researchers have determined that the occurrence of such a transition is influenced by the properties of the dataset (Chan et al., 2022a). Li et al. (2022); Lu et al. (2023) go further to associate the emergent capabilities seen during the learning with those that appear during the scaling of the model. Our research, however, is not centered on this transition phenomenon. We aim to delve into the learning plateaus phenomenon, which serves as a vital complement to understanding the intricacies of the transition phenomenon.
Understanding of the mechanism of in-context learning  Considerable research has been devoted to this vital topic, with most prior studies focusing on understanding the mechanism of in-context learning from the perspective of algorithm implementation. For example, a number of recent papers (von Oswald et al., 2022; Dai et al., 2023; Akyürek et al., 2022) have described in-context learning as akin to performing gradient descent. Additional research (Li et al., 2023; Bai et al., 2023) has interpreted in-context learning in terms of algorithm implementation and choice. Our study, however, takes a novel approach by examining the in-context learning mechanism through the lens of the weights and context components, offering a distinct perspective from earlier works.
@@ 3 Experimental Design
This paper employs a synthetic task to investigate the fundamental mechanisms behind learning plateaus. The use of a synthetic task is due to the fact that the intricacies of real tasks pose challenges in monitoring and comprehending the precise factors that influence the emergence of learning plateaus.
@@ 3.1 Dataset Construction
We propose a task using the Shapes3D (Kim & Mnih, 2018) dataset for a more controllable study. The experimental setting is shown in Fig. 2. Specifically, given a sequence of image and label pairs as context, the task involves predicting the label of the prompt image. Each image contains six different factors: object color, object shape, object scale, background color, floor color, and pose. We denote the factor as $e$ and the factor value of factor $e$ as $v^{(e)}$. For each sequence, we randomly choose a factor to generate the labels of the images, referring to this factor as the hidden factor $e_{h}$ for this sequence. For the two context sequences in Fig. 2, the hidden factor of Seq #1 is object color, and the correct label for the prompt image is 1 (object color is green). In Seq #2, for the same prompt image, the correct label is 3 (the object shape is a cube).
@@ 3.2 Analysis
To successfully tackle the in-context learning task, the network is required to discern the values of the six factors present in the prompt image, which relate to the weights component, as well as accurately determine the appropriate hidden factor for output based on the given contexts, which pertains to the context component. We break down the distribution ${\mathbb{P}}(y_{p}|{\mathbf{x}}_{p},s_{c})$ as follows.
@@
$\sum_{v_{p},m,e_{h}}\underbrace{{\mathbb{P}}(y_{p}|v_{p},m,e_{h})}_{\text{Properties of Task}}\underbrace{{\mathbb{P}}(v_{p}|{\mathbf{x}}_{p})}_{\text{weights}}\underbrace{{\mathbb{P}}(e_{h}|s_{c},m){\mathbb{P}}(m|s_{c})}_{\text{context}},$
@@ 3.3 Evaluation framework
In practice, since we cannot find the specific form of $g_{weights}$ and $g_{context}$, such that $f_{{\mathbf{w}},s_{c}}(x_{p})=g_{comb}(g_{weights}({\mathbf{x}}_{p}),g_{context}(s_{c}))$. We leverage the probe method to measure the goodness of the components in the inner representation of the Transformer as defined in Definition 3.3. We choose the layers in the Transformer that can produce the best probe results. We give the details of the probing framework in the Appendix A.3.
Probing methods and metrics  We use three metrics here. weights comp. score: accuracy of the probe model to predict $v_{{\mathbf{x}}}$ given ${\mathbf{x}}$, where “comp.” is short for component. context comp. score: the accuracy of the probe model to predict $e_{h}$ given $s_{c},{\mathbf{x}}_{p}$. Accuracy: We evaluate the prediction accuracy of the prompt example as the metric to evaluate the in-context learning performance. We give 39 context examples when evaluate context comp. score and accuracy.
All the results given in this paper are evaluated in the test set by default, as the performance of the test set is more aligned with the model’s performance on real situations.
@@ 3.4 Model, training and dataset detail
where $s^{(j)}\triangleq\{({\mathbf{x}}_{1},y_{1}),\cdots,({\mathbf{x}}_{j},y_{j})\}$, $\bm{l}$ denotes the loss function. ${\mathbf{x}}$ will be tokenized by VAE (Kingma & Welling, 2013) before being passed to Transformer. Note that the weights of VAE are pretrained and fixed during the whole experiments. The training loss in the dataset $S$, which contains $n$ sequence, is calculated as the average of loss over all training sequences, i.e.,
@@
We leave the details of the model, dataset, training design and configuration in the Appendix A.
@@ 4.1 Controlling of complexity
We aim to understand when the learning plateaus will happen by controlling the complexity of the tasks. Therefore, we will present the method for controlling the complexity in this part. Recall from Definition 3.1 that the data is generated following the formula ${\mathbb{P}}({\mathbf{x}},y,s_{c})=\sum_{m,e_{h}}{\mathbb{P}}(m){\mathbb{P}}(e_{h}){\mathbb{P}}({\mathbf{x}},y,s_{c}|m,e_{h})$. By altering the probabilities ${\mathbb{P}}(m)$, we can manipulate the resulting dataset. We first give two baseline configurations:
@@
When considering two distinct data configurations $D_{1}$ and $D_{2}$, the notation $D_{1}\Rightarrow D_{2}$ denotes the performance evaluation of a model on data setting $D_{2}$ after it has been trained on data setting $D_{1}$. Unless specified otherwise, we assume that the model is both trained and tested on the same data configuration.
@@ 4.2 Learning plateaus of task with different complexity
Learning Plateuas and Transition Point  The primary issue at hand involves learning plateaus and transition patterns. To begin with, we investigate the capability of the synthetic task to mimic the specific pattern depicted in Fig. 1B. Our scrutiny is directed toward the learning trajectory of the synthetic task. As demonstrated in Fig. 3A, it becomes evident that the synthetic task is successful in replicating the plateaus and transition pattern observed in the actual task.
The length of learning plateaus increases when increases the complexity of task  We delved into the relationship between task complexity and the duration of learning plateaus. We employed the entropy of the mapping function $m$’s distribution as an indicator of task complexity. To pinpoint the transition process, we identified the first epoch at which the model achieved a test accuracy greater than 0.17. This threshold was selected because the model’s accuracy remains below 0.17 before reaching the transition process and rises above this thereafter. As anticipated, more complex tasks necessitate longer learning plateaus (Fig. 3B). However, the relationship between the length of the plateaus and the entropy is not linear. With every unit increase in entropy, the extension of the learning plateau is marginal when the entropy is either low or relatively high; conversely, the growth in plateau length is more pronounced at intermediate levels of entropy.
Figure 4: Weights component and learning plateaus. A. The weights component score is increasing under $D_{\text{fix}}\Rightarrow D_{\text{fix}}$, while the weights component score is descreasing under the $D_{\text{rnd}}\Rightarrow D_{\text{rnd}}$ setting. Note that the “Weights” and “Context” in the figure are short for weights comp. score and context comp. score respectively. B. The weights component score after 50 epoch training decreases when increasing the complexity of the task. The dashed green line indicates the weights component score at the initialization point. C. The weights component score at 50 epoch negative correlates with the length of learning plateaus. The dashed green line indicates the weights component score at the initialization point.
4.3 Dysfunction of weights component